THE INTERACTIONS OF THE PEPTIDE FUNCTIONAL GROUP WITH WATER: A JET COOLED RO-TATIONAL STUDY OF THE FORMAMIDE-(H_{2}O)_{n} AND 2-AZETIDINONE-(H_{2}O)_{n} (n=1,2) CLUSTERS

\underline{JUAN C. LÓPEZ}, SUSANA BLANCO, RAQUEL SÁNCHEZ, ALBERTO LESARRI, and JOSÉ L. ALONSO, Grupo de Espectroscopía Molecular (GEM), Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47005 Valladolid, Spain.

Microsolvated clusters constitute models that allow the investigation of molecular properties in the transition from isolation to solvation and the understanding of solute-solvent interactions and solvent-enhanced chemical reactivity. In this work we have analyzed the jet-cooled rotational spectra of the formamide-(H_{2}O)_{n} and 2-azetidinone-(H_{2}O)_{n} (n=1, 2) clusters in order to investigate the diverse hydrogen bond interactions of water with the peptide functional group. Formamide (NH_{2}–CHO) and 2-azetidinone (C_{3}H_{5}NO), a lactam four-membered ring that forms part of penicillin- and cephallosporintype antibiotics, are two of the most simple molecules carrying a peptide group. A molecular beam Fourier Transform microwave spectrometer^{a} equipped with a heating nozzle^{b} has been used for this research. We have observed thre 1:1 forms of the formamide-H_{2}O complex, two 1:1 forms of the 2-azetidinone-H_{2}O complex and the trimers formamide-(H_{2}O)_{2} and 2-azetidinone-(H_{2}O)_{2}. The most stable conformer of the formamide-H_{2}O complex, previously observed by Lovas et al.,^{c} and the most stable form of the 2-azetidinone-H_{2}O complex are stabilized by two hydrogen bonds O-H⋯O=C and N-H⋯O with water closing a cycle with the peptide functional group. In the second most stable 1:1 conformers water is placed on the other side of the carbonyl group stabilized by O-H⋯O=C hydrogen bonds and by C-H⋯O weak hydrogen bonds. In the third formamide-H_{2}O form, water is bonded to the amino group by a N-H⋯O hydrogen bond. For the trimers, the two molecules of water close a cycle with the peptide group with three different hydrogen bonds (O-H⋯O=C, O-H⋯O and N-H⋯O). The spectra of the parent and several D/H, ^{15}N/^{14}N, ^{13}C/^{12}C and ^{18}O/^{16}O isotopomers have been measured in order to investigate the structures of these clusters.

a J. L. Alonso, F. Lorenzo, J. C. López, A. Lesarri, S. Mata and H. Dreizler; Chem. Phys., 218, 267 (1997)

b S. Blanco, J. C. López, J. L. Alonso, P. Ottaviani and W. Caminati; J. Chem. Phys., 119, 880 (2003)

c F. J. Lovas, R. D. Suenram, G. T. Fraser, C. W. Gillies and J. Zozom, J. Chem. Phys., 88 722 (1988)